OpenStudy (anonymous):

Solve for x: [2x-7]=13

OpenStudy (anonymous):

x=10

OpenStudy (anonymous):

you need to solve: \[2x-7\ge 13, 2x-7<14\] assuming those brackets are the greatest integer function.

OpenStudy (anonymous):

transpose -7 to right hand side 2x- 7 = 13 2x = 13 + 7 2x = 20 divide both sides by 2 2x/2 = 20/2 x = 10

OpenStudy (anonymous):

what are those brackets? are you treating them just like parenthesis? The greatest integer function gives the greatest integer less than or equal to x, for example: [5.5] = 5 [4.999999999] = 4 [-3.2] = -4